OpenStudy (anonymous):

The slope of the line passing through P(1,4) and Q(x,1) is -3/5. What is the value of x? a. 6 b. 7 c. 9 d. 10

OpenStudy (anonymous):

c

OpenStudy (decentnabeel):

So for every 3 you go forwards in X, Y drops by 5. Easiest way is to find the equation, imho. So -3/5x + z=Y z = unknown so we have 1 pt to use (P). so -3/5 x 1 (our x)= -3/5 but the Y value is 4, so the difference is 4 - (-3/5) = 4.6 so the equation is -3/5x + 4.6 = Y so if Y=1 for the second pt, Q, plug it into the equation to get 1= -3/5x + 4.6 rearrange to get -3.6=-3/5x divide both sides by -3/5 x=6

OpenStudy (decentnabeel):

no @slade answer is a

OpenStudy (anonymous):

i didnt answer this Q

OpenStudy (anonymous):

i just got on

OpenStudy (decentnabeel):

so expaln it

OpenStudy (anonymous):

i gonna go change my password

OpenStudy (decentnabeel):

are you undersatnd @Gabrielle_Life

OpenStudy (decentnabeel):

why the answer is c @slade

OpenStudy (lynfran):

@DecentNabeel your work is correct

OpenStudy (anonymous):

yeah i am so confused

OpenStudy (decentnabeel):

thank you @LynFran

OpenStudy (decentnabeel):

why confused expaln it @Gabrielle_Life

OpenStudy (anonymous):

never mind i got it

OpenStudy (decentnabeel):

ok @Gabrielle_Life